# Multivariable Painlevé-II equation: connection formulas for asymptotic solutions

### 23 Mar 2026 v1

@@ V Excitations after unstable vacuum decay
${\cal H}(t)=\frac{{\bm{P}}^{2}}{2}-t\frac{{\bm{X}}^{2}}{2}+\frac{{\bm{X}}^{4}}{2}+\frac{eu_{2}^{2}}{2},$ (48)
@@
${\cal J}_{1,2}\equiv\frac{\alpha_{1,2}^{2}}{2}.$
Similarly, $I_{1}$ and $I_{2}$ from Eq. (30) are identified as asymptotically conserved adiabatic invariants as $t\rightarrow+\infty$. Then, $\varphi_{1,2}$ and $\phi_{1,2}$ are the angles canonically conjugated to, respectively, ${\cal J}_{1,2}$ and $I_{1,2}$.
According to [31, 30], this classical dynamics appears as a saddle-point solution for a quantum field theory describing a time-dependent passage through a second-order quantum phase transition. The difference between ${\cal J}_{1,2}$ and $I_{1,2}$ captures inevitable nonadiabatic effects as the classical phase space trajectory passes through a separatrix. Physically, this event corresponds to crossing the critical point of a phase transition in the corresponding field theory [30]. Semiclassically, the actions are related to the numbers of produced quasiparticles: $I_{1}/\hbar$ for the number of Higgs bosons and $I_{2}/\hbar$ for the number of Goldstone bosons. The latter acquire a small mass due to the symmetry breaking term $\propto e$ in Eq. (48). The analysis in [31] did not provide explicit connection formulas but made several observations that can now be verified.
Thus, a useful feature of Eqs. (33) and (35) is that $I_{1}$ and $I_{2}$ depend on $e$ only through the angles $\Phi_{1}$ and $\Phi_{2}$, which themselves depend linearly on the initial phases $\varphi_{1}$ and $\varphi_{2}$. Therefore, $\langle I_{1}\rangle$ and $\langle I_{2}\rangle$ are independent of $e$, where the brackets denote averaging over a uniform distribution of $\varphi_{1}$ and $\varphi_{2}$. This invariance of the averaged actions was conjectured in [31] based on less rigorous analytical arguments and numerical simulations, and is now confirmed.
Next, the solution in Eqs. (32)-(36) exhibits asymmetry amplification, which was the main finding in [31]. No matter how small the symmetry breaking parameter $e$ is, the dynamics for $u_{1}(t)$ and $u_{2}(t)$ strongly differ from one another as $t\rightarrow+\infty$. While $u_{1}(t)$ has a monotonically growing component, the asymptotic solution for $u_{2}(t)$ oscillates around zero. Note that if $e$ changes sign, the asymptotic solutions in Eqs. (III.1) and (29) are obtained by switching indices $1$ and $2$. The solution at $e=0$ reduces to the standard P-II solution, which is unstable in the present context.
@@
$\langle I_{2}\rangle+2\langle I_{1}\rangle\approx\frac{1}{2\pi}\ln\left(\frac{1}{2\pi\cal J}\right).$ (50)
@@
$\langle I_{1}\rangle\approx\frac{1}{4\pi}\left[\ln\left(\frac{1}{2\pi{\cal J}}\right)-c_{1}\right],$ (51)
@@
Expressions for $\langle I_{1}\rangle$ and $\langle I_{2}\rangle$ in Eqs. (51) and (53) were conjectured in [31] for ${\cal J}\ll 1$ without an analytical derivation of $c_{1,2}$, while numerical simulations in [31] estimated that $c_{1}\approx 1.14$ and $c_{2}\approx 1.19$, with an uncertainty in the last significant digit. This agrees well with the analytically derived values of these coefficients in Eq. (53). Curiously, the value of the double phase integral in Eq. (52), $\approx 1.166$, is the famous square-lattice spanning tree constant [32].

### 27 Mar 2026 v2

@@ V Excitations after unstable vacuum decay
${\cal H}(t)=\frac{{\bm{P}}^{2}}{2}-t\frac{{\bm{X}}^{2}}{2}+\frac{{\bm{X}}^{4}}{2}+\frac{\varepsilon u_{2}^{2}}{2},$ (48)
@@
${\cal I}_{1,2}\equiv\frac{\alpha_{1,2}^{2}}{2}.$
Similarly, $I_{1}$ and $I_{2}$ from Eq. (30) are identified as asymptotically conserved adiabatic invariants as $t\rightarrow+\infty$. Then, $\varphi_{1,2}$ and $\phi_{1,2}$ are the angles canonically conjugated to, respectively, ${\cal I}_{1,2}$ and $I_{1,2}$.
According to [32, 31], this classical dynamics appears as a saddle-point solution for a quantum field theory describing a time-dependent passage through a second-order quantum phase transition. The difference between ${\cal I}_{1,2}$ and $I_{1,2}$ captures inevitable nonadiabatic effects as the classical phase space trajectory passes through a separatrix. Physically, this event corresponds to crossing the critical point of a phase transition in the corresponding field theory [31]. Semiclassically, the actions are related to the numbers of produced quasiparticles: $I_{1}/\hbar$ for the number of Higgs bosons and $I_{2}/\hbar$ for the number of Goldstone bosons. The latter acquire a small mass due to the symmetry breaking term $\propto\varepsilon$ in Eq. (48). The analysis in [32] did not provide explicit connection formulas but made several observations that can now be verified.
Thus, a useful feature of Eqs. (33) and (35) is that $I_{1}$ and $I_{2}$ depend on $\varepsilon$ only through the angles $\Phi_{1}$ and $\Phi_{2}$, which themselves depend linearly on the initial phases $\varphi_{1}$ and $\varphi_{2}$. Therefore, $\langle I_{1}\rangle$ and $\langle I_{2}\rangle$ are independent of $\varepsilon$, where the brackets denote averaging over a uniform distribution of $\varphi_{1}$ and $\varphi_{2}$. This invariance of the averaged actions was conjectured in [32] based on less rigorous analytical arguments and numerical simulations, and is now confirmed.
Next, the solution in Eqs. (32)-(36) exhibits asymmetry amplification, which was the main finding in [32]. No matter how small the symmetry breaking parameter $\varepsilon$ is, the dynamics for $u_{1}(t)$ and $u_{2}(t)$ strongly differ from one another as $t\rightarrow+\infty$. While $u_{1}(t)$ has a monotonically growing component, the asymptotic solution for $u_{2}(t)$ oscillates around zero. Note that if $\varepsilon$ changes sign, the asymptotic solutions in Eqs. (III.1) and (29) are obtained by switching indices $1$ and $2$. The solution at $\varepsilon=0$ reduces to the standard P-II solution, which is unstable in the present context.
@@
$\langle I_{2}\rangle+2\langle I_{1}\rangle\approx\frac{1}{2\pi}\ln\left(\frac{1}{2\pi\cal I}\right).$ (50)
@@
$\langle I_{1}\rangle\approx\frac{1}{4\pi}\left[\ln\left(\frac{1}{2\pi{\cal I}}\right)-c_{1}\right],$ (51)
@@
Expressions for $\langle I_{1}\rangle$ and $\langle I_{2}\rangle$ in Eqs. (51) and (53) were conjectured in [32] for ${\cal I}\ll 1$ without an analytical derivation of $c_{1,2}$, while numerical simulations in [32] estimated that $c_{1}\approx 1.14$ and $c_{2}\approx 1.19$, with an uncertainty in the last significant digit. This agrees well with the analytically derived values of these coefficients in Eq. (53). Curiously, the value of the double phase integral in Eq. (52), $\approx 1.166$, is the famous square-lattice spanning tree constant [33].